Fermi National Laboratory


Accelerator Update

Monday February 23
The day shift began with Operations monitoring store 3252, with MiniBooNE taking beam, and with the Antiproton Source (Pbar) stacking.

At 9:49 AM, a Pbar expert began a Debuncher study.

A TeV expert began a collimator and separator study at 10:59 AM. The study lasted for one and a half hours.

The Main Injector's RF sector suffered a vacuum burst at 11:47 AM. The vacuum pumps all reset okay.

A Switchyard (SY) expert finished some target scans at 12:35 PM.

At 1:47 PM, a Booster power supply's output went to zero and caused beam problems. Operators found the troubled supply and managed to get it operating again.

A Tevatron (TeV) expert began an end-of-store study at 3 PM.

At 4:11 PM, Operations terminated store 3252.

The Neutron Therapy Facility (NTF), which now only does studies, began taking beam at 4:51 PM.

Operations put the TeV into shot setup at 5:20 PM. At 7:07 PM, Operations established store 3254 with an initial luminosity of 51.39E30. Pbar resumed stacking.

At 7:43 PM, Pbar lost its stack. Operators discovered an Accumulator power supply (A:IB) not regulating properly. Experts were notified and operators called in two power supply technicians. At 11:38 PM, experts reported that they were still working on the power supply.

Tuesday February 24
The midnight shift began with Operations monitoring store 3254, with MiniBooNE taking beam, and with power supply repairs in Pbar.

At 1:36 AM, a TeV sector B1 wet engine alarmed on over-temperature. Cryo system techs from CHL investigated and said it was an electrical problem.

The Pbar power supply experts had A:IB up and running by 2:40 AM. Pbar resumed stacking.

At 4:48 AM, a Booster power supply (GMPS) tripped off for the first of five times in one and a half hours.

Operations reported at 5:42 AM, that stacking had stopped due to some Pbar beam problems. Operators called an expert and worked on tuning the beam for about an hour.

At 7:13 AM, Pbar's A:IB stopped regulating again. Operators contacted experts and then tried various repairs. With the help of the Duty Electrician, operators managed to get the supply to regulate properly.

At 7:17 AM, Booster went down again due to a power supply problem (CEPNE1). This problem, of course, stopped all downstream beam. Experts investigated.

A Recycler tech took down a MI power supply briefly, at 8:31 AM, in order to work on some Recycler equipment.

A Booster expert reported that they were going to try and run Booster beam at 8:34 AM.

Pbar resumed stacking at 8:51 AM.

At 8:52 AM, Switchyard and MI experts began a SY study. Experts established beam to the SY dump at 9:32 AM. Under expert supervision, Operations began sending beam, at 10 AM, to the Meson Test experiment at MT6.

At 3:40 PM, beam in the 400 MeV area of Linac tripped off for the first of five times in four minutes. After the last trip operators call an expert.

SY experts completed their studies at 7:26 PM.

Operations reported at 11:57 PM that they had been having problems with a magnet (MV1), which was reducing MiniBooNE's beam intensity.

Wednesday February 25
The midnight shift began with Operations monitoring store 3254, with Pbar stacking, and with MiniBooNE taking beam.

At 6:04 AM, a TeV expert began an end-of-store study.

Today's Plans
The plans for today include terminating the store as soon as the end-of-store studies get completed. Then time will be given for cryo techs to work on a wet engine and to allow experimenters at the Collider Detector at Fermilab and D-zero to access their collision halls. Later in the morning Operations will put the TeV into shot setup.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 2/25/2004   email Fermilab